How do I find a missin angle in a heptagon? I am given angles 142 130 122 105 and 110
Answer by stanbon(75887)

How do I find a missing angle in a heptagon? I am given angles 142 130 122 105 and 110
------------
The sum of the interior angles of an n-gon is (n-2)180
---------------------------------------
You've listed 5 angles so I guess you are looking for the 6th.
--------------------
(6-2)180 = 4*180 = 720
---------------------------
720-(142+130+122+105+110) = your answer
===========================================
Cheers,
Stan H.